40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11021     Entry time: Fri Feb 13 03:44:56 2015
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Held using ALS for a while at "0" CARM offset with PRMI 

[Jenne, Rana]

We wanted to jump right in and see if we were ready to try the new "ALS fool" loop decoupling scheme, so we spent some time with CARM and DARM at "0" offset, held on ALS, with PRMI locked on REFL33I&Q (no offsets).  Spoiler alert:  we weren't ready for the jump.

The REFL11 and AS55 PDs had 0dB analog whitening, which means that we weren't well-matching our noise levels between the PD noise and the ADC noise.  The photodiodes have something of the order nanovolt level noise, while the ADC has something of the order microvolt level noise.  So, we expect to need an analog gain of 1000 somewhere, to make these match up.  Anyhow, we have set both REFL11 and AS55 to 18dB gain. 

On a related note, it seems not so great for the POX and POY ADC channels to be constantly saturated when we have some recycling gain, so we turned their analog gains down from 45dB to 0dB.  After we finished with full IFO locking, they were returned to their nominal 45dB levels. 

We also checked the REFL33 demod phase at a variety of CARM offsets, and we see that perhaps it changes by one or two degrees for optimal rotation, but it's not changing drastically.  So, we can set the REFL33 demod phase at large CARM offset, and trust it at small CARM offset.

We then had a look at the transmon QPD inputs (before the dewhitening) for each quadrant.  They are super-duper saturating, which is not so excellent. 
QPDsaturation_12Feb2015.pdf

We think that we want to undo the permanently-on whitening situation.  We want to make the second stage of whitening back to being switchable.  This means taking out the little u-shaped wires that are pulling the logic input of the switches to ground.  We think that we should be okay with one always on, and one switchable.  After the modification, we must check to make sure that the switching behaves as expected.  Also, I need to figure out what the current situation is for the end QPDs, and make sure that the DCC document tree matches reality.  In particular, the Yend DCC leaf doesn't include the gain changes, and the Xend leaf which does show those changes has the wrong value for the gain resistor.

After this, we started re-looking at the single arm cancellation, as Rana elogged about separately.

ALSfool_12Feb2015.pdf

Attachment 1: QPDsaturation_12Feb2015.pdf  142 kB  | Hide | Hide all | Show all
QPDsaturation_12Feb2015.pdf
Attachment 2: ALSfool_12Feb2015.pdf  12 kB  | Show | Hide all | Show all
ELOG V3.1.3-